OpenStudy (anonymous):

find the quotient 7 1/8 divided by 4 4/5

OpenStudy (anonymous):

simplfy answer

OpenStudy (anonymous):

what does 7 1/8 equal to?

OpenStudy (anonymous):

\[a \frac{ b }{ c }=\frac{ a*c+b }{ c }\]

OpenStudy (anonymous):

you there, or should i go help someone else?

OpenStudy (anonymous):

no im doing it

OpenStudy (anonymous):

57/8

OpenStudy (anonymous):

4 4/5=

OpenStudy (anonymous):

17 over 5

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

4*4+1

OpenStudy (anonymous):

where did you get these numbers from?

OpenStudy (anonymous):

i did it wrong

OpenStudy (anonymous):

ok whats the answer now

OpenStudy (anonymous):

4*5+4 = 24 over 5

OpenStudy (anonymous):

\[\frac{ \frac{ 57 }{ 8 } }{ \frac{ 24 }{ 5 } }\] when dividing fraction by fraction it is same as _________?

OpenStudy (anonymous):

timesing them

OpenStudy (anonymous):

times them the way they are?

OpenStudy (anonymous):

8*57

OpenStudy (anonymous):

no

OpenStudy (anonymous):

57*24?

OpenStudy (anonymous):

\[\frac{ \frac{ a }{ b } }{ \frac{ c }{ d } }=\frac{ a }{ b }*\frac{ d }{ c }\]

OpenStudy (anonymous):

456 over 120

OpenStudy (anonymous):

how did you get 456

OpenStudy (anonymous):

57 times8

OpenStudy (anonymous):

no it will be 57/8*5/24

OpenStudy (anonymous):

i need it in fraction form

OpenStudy (anonymous):

that gives me 1.484375

OpenStudy (anonymous):

\[\frac{57 }{ 8 }*\frac{ 5 }{ 24 }=\frac{ ? }{ ? }\]

OpenStudy (anonymous):

285 over192

OpenStudy (anonymous):

what can you divide both numbers with

OpenStudy (anonymous):

192.

OpenStudy (anonymous):

no to simplify the answer there is a common number you can divide both numbers with Eample \[\frac{2}{4}=\frac{1}{2}\]

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

\[\frac{ 2 }{ 4 }=\frac{ 2/2 }{ 4/2 }=\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

what do i do this for

OpenStudy (anonymous):

divide both numbers by 3

OpenStudy (anonymous):

152 over 40

OpenStudy (anonymous):

no 285/3=?

OpenStudy (anonymous):

95

OpenStudy (anonymous):

192/3?

OpenStudy (anonymous):

64

OpenStudy (anonymous):

so whats the answer?

OpenStudy (anonymous):

95 over 40

OpenStudy (anonymous):

where did you get 40 from?

OpenStudy (anonymous):

95 over 64 i mean

OpenStudy (anonymous):

correct